OpenStudy (anonymous):

how do i graph sec and csc on my calculator. it come up all wrong when i try

OpenStudy (anonymous):

What kind of calculator is it?

OpenStudy (anonymous):

they are funky graphs

OpenStudy (anonymous):

ti 83 plus. and the problem is 2sec(8(x+1)

OpenStudy (anonymous):

i have tried to type it in as. 2(1/cos(8(x+1))

OpenStudy (anonymous):

Hmm, that should have worked.

OpenStudy (anonymous):

are there 3 ))) at the end of that? or did you type it as is?

OpenStudy (anonymous):

i agree.... can you explain how to graph nonsinusotal graphs by hang?

OpenStudy (anonymous):

@Euler271 what do you mean?

OpenStudy (anonymous):

on both times that you wrote what you tried to type in your calculator you forgot a ')' [closed bracket]

OpenStudy (anonymous):

oh i can try that

OpenStudy (anonymous):

Try adjusting your window parameters.

OpenStudy (anonymous):

its still not working, the original problem was h(x)=2sec(pi/4(x+1)), and asked to find the period and horizontal shift for the problem. would i use pi/4 instead?

OpenStudy (anonymous):

Change the Xmin to \(-\dfrac{\pi}{8}\), Xmax to \(\dfrac{\pi}{8}\), Ymin to -5, and Ymax to 5.

OpenStudy (anonymous):

okay, its not graphing like normal, thank you. but in my book it says x=-1 y=2 and x=0 y=3 and this is not showing

OpenStudy (anonymous):

That last point must be typo then. \[y=2\sec\left(\frac{\pi}{4}(x+1)\right)~~\Rightarrow~~x=0,~y=\frac{2}{\cos{\frac{\pi}{4}}}=\frac{2}{\frac{1}{\sqrt2}}=2\sqrt 2\not=3\]

OpenStudy (anonymous):

okay, thank you alot. how would i graph these by hand?

OpenStudy (anonymous):

like would i graph the cos of this and to use as a guide, and how would i find the asymototes?

OpenStudy (anonymous):

When graphing trig functions by hand, I would plug in easy values of \(x\). By easy, I mean values that make it easy to find \(y\). For example, given \(y=\sin 2x\), an "easy" \(x\) would be \(x=\dfrac{\pi}{4}\), which gives you \(y=\sin2\left(\dfrac{\pi}{4}\right)=\sin\dfrac{\pi}{2}=1\).

OpenStudy (anonymous):

i see, how aobut like y=tan(x) and y=csc(x)?